the units of a two-digit numeral is 5 more than the tens digit the number is 3 times the sum of its digits.find the numeral
Found 2 solutions by dturpin53, solver91311:
Answer by dturpin53(1)

let a be the 10's digit and let b be the units digit.
From the stated problem the units digit is 5 more than the tne's digit or in equation form:
b = a + 5
The number is 3 times the sum of the digits or in equation form we have:
10a + b= 3( a + b)= 3a + 3b
substituting the b value from the first equation into the second we have:
10a + a + 5 = 3a + 3(a +5)= 3a + 3a + 15
11a + 5 = 6a + 15
5a = 10
a = 2
b = a + 5 = 2 + 5 = 7
is this correct, the number is 27, and the sum of digits is 9 and 3*9 = 27

Answer by solver91311(24713)



Let represent the 10s digit. Then must represent the units digit.

The number is then , and the sum of the digits is , and three times the sum of the digits is then , so:



Solve for (which you know has to be 4 or less -- otherwise adding 5 to get the units digit would result in something other than a digit)


John